A linear demand curve is given by Q_D=225-12P
Graph the demand curve. Be sure to clearly indicate the vertical intercept.
Now assume the price of good is $10, calculate the following
The quantity demanded
The amount paid
The consumer surplus
The total willingness to pay
On your graph in part 1, clearly illustrate part 2 (i.e. indicate the quantity demanded, amount paid, consumer surplus and total wiliness to pay
1
Expert's answer
2020-10-25T18:25:05-0400

If the price of good is $10, then the quantity demanded is:

QD = 225 - 12×10 = 105 units.

The amount paid is:

10×105 = $1,050.

The consumer surplus is:

CS = 0.5×105×(225/12 - 10) = $459.38.

The total willingness to pay is: 1,050 + 459.38 = $1,509.38.